# Late Jets, Early Sparks: Illuminating the Pre-Maximum Bumps in Superluminous Supernovae

### v1

@@ Abstract
Superluminous supernovae (SLSNe) radiate $\gtrsim 10-100$ times more energy than ordinary stellar explosions, implicating a novel power source behind these enigmatic events. One frequently discussed source, particularly for hydrogen-poor (Type I) SLSNe, is a central engine such as a millisecond magnetar or accreting black hole. Both black hole and magnetar engines are expected to channel a fraction of their luminosity into a collimated relativistic jet. Using 3D relativistic hydrodynamical simulations, we explore the interaction of a relativistic jet, endowed with a luminosity $L_{\rm j}\approx 10^{45}\,{\rm erg\,s^{-1}}$ and duration $t_{\rm eng}\approx$ days compatible with those needed to power SLSNe, launched into the envelope of the exploding star. The jet successfully breaks through the expanding ejecta and its shocked cocoon powers ultraviolet/optical emission lasting several days after the explosion and reaching a peak luminosity $\gtrsim 10^{44}$ erg s-1, corresponding to a sizable fraction of $L_{\rm j}$. This high radiative efficiency is the result of the modest adiabatic losses the cocoon experiences due to the low optical depths of the enlarged ejecta at these late times, e.g., compared to the more compact stars in gamma-ray bursts. The luminosity and temperature of the cocoon emission match those of the “bumps” in SLSNe light curves observed weeks prior to the optical maximum in many SLSNe. Confirmation of jet breakout signatures by future observations (e.g., days to weeks-long internal X-ray emission from the jet for on-axis observers, spectroscopy confirming large photosphere velocities $v/c\gtrsim 0.1$, or detection of a radio afterglow) would offer strong evidence for central engines powering SLSNe.
@@ 1 Introduction
The collapse of a rotating massive star giving birth to a magnetized spinning compact object -- either a Kerr black hole or a neutron star -- offers a compelling scenario for generating some of the most energetic transients in the Universe, including gamma-ray bursts (GRB) and superluminous supernovae (SLSNe). While most GRBs with long durations are accompanied by stripped-envelope supernovae with abnormally large kinetic energies (“broad-lined” Type Ibc SNe; e.g., Stanek et al. 2003; Woosley & Bloom 2006; Liu & Modjaz 2016), SLSNe are instead distinguished by their large radiated energies, which can exceed those of ordinary stellar explosions by a factor of $\gtrsim 10-100$ (e.g., Gal-Yam 2018; Nicholl 2021 for reviews). Type I (i.e., hydrogen-poor) SLSNe, in particular, may be associated with the birth of a millisecond (ms) neutron star with a strong magnetic field (“ms magnetar”; e.g., Kasen & Bildsten 2010; Woosley 2010; Nicholl et al. 2017; Vurm & Metzger 2021). The magnetized wind which accompanies the first tens of seconds in the life of a ms magnetar can feed a relativistic jet with the necessary luminosity and baryon loading (Lorentz factor) to power a GRB (e.g., Wheeler et al. 2000; Thompson et al. 2004; Bucciantini et al. 2008; Metzger et al. 2011), though the accretion-powered jet of a Kerr black hole remains a more commonly invoked scenario (e.g., Woosley 1993; MacFadyen & Woosley 1999; Gottlieb et al. 2023).
@@
What distinguishes the central engines of GRBs versus those that power SLSNe? Although the total energetics of both systems can be comparable, their main distinction is the duration of the central engine activity (Metzger et al., 2015). While GRB engines typically emit a sizable fraction of their energy (Panaitescu & Kumar, 2002) over seconds to minutes (comparable to the duration of the gamma-ray emission; e.g., Bromberg et al. 2012), the engines of SLSNe must instead heat the supernova ejecta continuously from within over much longer timescales of days to weeks. This heating being substantially delayed after the time of the stellar explosion is crucial to explaining the high radiative output of SLSNe. A delay ensures that less deposited energy is “wasted” through adiabatic losses by occurring at later times after the optical depth through the expanding ejecta has dropped. Fig. 1 shows estimates of the engine duration and luminosity required to explain the rise times and peak luminosities of a large sample of SLSNe-I (Gomez et al., 2024).
@@
In magnetar scenarios, the engine duration is typically related to the magnetic dipole spin-down timescale $t_{\rm sd}\propto B^{2}$ which varies considerably with the surface magnetic field strength ($B\sim 10^{14}$ G to power SLSNe vs. $B\gtrsim 3\times 10^{15}$ G to power a typical long GRB). Even at fixed $B$, differences in the magnetar’s mass accretion rate (Metzger et al., 2018) or the inclination angle of the magnetic dipole relative to the spin axis (Margalit et al., 2018) can lead to large differences in the jet power. So-called “ultra-long” GRBs (e.g., Levan et al. 2014) require engine durations of hours, commensurate with their longer gamma-ray emission; these events may represent an intermediate case in which the supernova’s luminosity is boosted only modestly above the baseline level set by 56Ni decay (Greiner et al., 2015; Metzger et al., 2015), the latter dominating in ordinary GRB supernovae (e.g., Woosley et al. 2002).
@@
While a variety of indirect evidence supports a compact object residing at the center of at least some SLSNe-I (e.g., Lunnan et al. 2015; Nicholl et al. 2019; Eftekhari et al. 2019; Vurm & Metzger 2021; Omand & Jerkstrand 2023), a central engine is not the only plausible power source for these events. A commonly invoked alternative is shock interaction between the fast supernova ejecta and a compact wind or other circumstellar medium (CSM) that surrounds the progenitor star at the time of explosion (e.g., Chevalier & Irwin 2011; Chatzopoulos et al. 2013; Lunnan et al. 2018; Fraser 2020; Aguilera-Dena et al. 2020; Stevance & Eldridge 2021). Given the challenges in cleanly distinguishing engine from CSM interaction models, alternative tests of the engine hypothesis are sorely needed. Gamma-ray emission from a magnetar nebula on timescales of months to years after the ejecta becomes optically-thin offers one compelling test (e.g., Murase et al. 2016; Vurm & Metzger 2021); however, the low predicted gamma-ray flux at these late times makes detection challenging in all but the closest explosions (Acharyya et al., 2023). A more promising test is one that probes the central engine at earlier times closer to its maximal power.
@@
If a rotating compact object powers both GRBs and SLSNe, it is natural to ask whether SLSNe should also be accompanied by ultra-relativistic jets. While the luminosities of SLSN engines are lower than GRB jets, the density of the external medium through which a putative SLSN jet must propagate is also lower. This is due to the larger ejecta radius $R_{\rm ej}\approx\,{v_{\rm ej}}t\sim 10^{13}-10^{14}$ cm at such late times tsimilar-$t\sim$ days--weeks after the explosion. Here, $\,{v_{\rm ej}}\approx 5000$ km s-1 is the speed of the supernova ejecta, which for typical stripped-star progenitor radii $R_{\star}\sim 10^{11}$ cm, achieves homologous expansion on a timescale $\gtrsim R_{\star}/\,{v_{\rm ej}}\sim 10^{2}$ s. Analytic estimates show that even jet luminosities $\,{L_{\rm j}}\simeq 10^{45}$ erg s-1, much lower than for GRBs but comparable to those required of many SLSNe-I engines (Fig. 1), are sufficient for jet breakout through the homologously expanding stellar ejecta (e.g., Quataert & Kasen 2012; Margalit et al. 2018; Gottlieb & Nakar 2022). In particular, for a jet of half-opening angle $\theta_{j}$ and luminosity $\,{L_{\rm j}}$, breakout can occur on a timescale (Eq. 45 in Gottlieb & Nakar, 2022):
$t_{\rm bo}\approx 1.5\,\frac{E_{\rm KE}}{10^{51}\,\,{{\rm erg}}}\frac{10^{45}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}}{\,{L_{\rm j}}}\left(\frac{\theta_{j}}{10^{\circ}}\right)^{4}\,\,{{\rm d}}\,,$ (1)
@@
where the prefactor assumes a non-magnetized jet propagation, as expected for extended-duration weak jets, and other quantities including the ejecta kinetic energy $E_{\rm KE}$ are normalized to characteristic values (Fig. 1). The jet breakout times of days indicated by Eq. (1) are notably similar to the typical engine durations (e.g., magnetar spin-down time) required to power the bulk of the SLSNe emission.
@@
Figure 1: Required engine peak luminosity $L_{\rm eng}$ and duration at this luminosity, $t_{\rm eng}$ (e.g., magnetar spin-down luminosity and timescale, respectively) to explain the observed peak luminosities, $L_{\rm pk}$, and rise times, $t_{\rm pk}$, of a large sample of SLSNe-I light curves (Gomez et al., 2024), for two different assumptions about the ejecta mass, $M_{\rm ej}=10M_{\odot}$ (black stars) and $3M_{\odot}$ (blue squares), bracketing the typical range inferred for Type I SLSNe (Nicholl et al., 2017). For comparison, we show the approximate range of timescales and luminosities of the early pre-maximum bumps observed in the light curves of SLSNe-I (shaded blue ellipse; see e.g. Nicholl & Smartt 2016). A multiday-lasting jet from the same central engine responsible for powering the bulk of the SLSNe is thus capable of powering the early bumps, assuming the jet luminosity, $\,{L_{\rm j}}$, is a fraction $\sim 1-10\%$ of $L_{\rm eng}$. The assumed jet properties for the numerical simulation presented in this paper are illustrated with an orange square, while left-right arrows indicate the degeneracy in engine properties $L_{\rm eng}\propto t_{\rm eng}^{-1}$ that result in the same hydrodynamic evolution for fixed ejecta properties. Details: For each event, we estimate the ejecta kinetic energy, $E_{\rm KE}=M_{\rm ej}\,{v_{\rm ej}}^{2}/2$, by equating the observed supernova rise-timescale to the diffusion timescale for a centrally-concentrated energy source, $t_{\rm diff}\simeq 0.4(\kappa M_{\rm ej}/\,{v_{\rm ej}}c)^{1/2}$ (Khatami & Kasen, 2019). Assuming the explosion energy $E_{\rm KE}$ comes exclusively from the engine, i.e. $E_{\rm eng}=L_{\rm eng}t_{\rm eng}=E_{\rm KE}$, we estimate the required engine timescale assuming the engine luminosity decays similar to magnetic-dipole spin-down $L_{\rm pk}=(E_{\rm eng}/t_{\rm eng})(1+t_{\rm pk}/t_{\rm eng})^{-2}$, for $t_{\rm eng}<t_{\rm pk}$. If the engine contributes only a fraction of the explosion energy $E_{\rm KE}$, then a greater engine duration is needed to explain the observed luminosities and points will shift towards the lower-right-hand part of the diagram as indicated by the arrow.
@@
Do observations rule out jets in SLSNe? With one interesting exception (Levan et al. 2013), no X-ray emission has been detected to faint limits from a sample of 26 SLSNe-I on timescales of months to years after the explosion (e.g., Margutti et al. 2018). However, this is not necessarily constraining because the spectral energy distribution of any (internal) jetted emission is uncertain; and even if the jet efficiently powers X-rays, they may only be detectable at early times and for observers within the narrow opening angle of the jet. In principle, radio afterglow emission from the jet as it interacts with surrounding CSM may be detectable at late times even for slightly off-axis observers; however, while again (with one or two interesting exceptions; Eftekhari et al. 2019; Margutti et al. 2023111In PTF10hgi (Eftekhari et al., 2019; Mondal et al., 2020), radio emission detected starting 7.5 years after the explosion may originate from a magnetar nebula (e.g., Margalit & Metzger 2018; Omand et al. 2018), while the late radio emission in SN 2017ens (Margutti et al., 2023) originates from the supernova ejecta colliding with CSM.) no radio detections of SLSNe-I have been made (e.g., Coppejans et al. 2018; Hatsukade et al. 2021; Eftekhari et al. 2021). As we shall discuss, the resulting constraints on the jet power are sensitive to the observer viewing angle, CSM density, and uncertain microphysical parameters of the shock.
@@
A cleaner signature of a successful jet in SLSNe-I may come at earlier times, on timescales comparable to the engine duration of days to weeks (Margalit et al., 2018). In ordinary GRB jets, the compact size of the shocked stellar debris (the so-called “cocoon”; e.g., Gottlieb et al. 2018a) results in most of the cocoon radiation being initially trapped and only subsequently released after many expansion times. By contrast, for the more sustained jet-ejecta interaction expected for longer-lived SLSNe jets interacting with substantially larger ejecta, the lower optical depths (smaller trapping radius in the ejecta frame) will alter the observed cocoon light curve, allowing it to more faithfully track energy input from the engine.
@@
A suggestive hint comes from the early maxima (“bumps”) of luminosity $\sim 10^{43}-10^{45}$ erg s-1 observed in some SLSNe-I optical light curves on timescales of days to a week after the explosion, well before the main supernova peak (Leloudas et al., 2012; Nicholl et al., 2015). Although such pre-maximum bumps are not detected in all events due to observational limitations, Nicholl & Smartt (2016) show that they are likely common features in SLSNe. Subsequent studies have shown that early maxima are not present in all SLSNe, and that their properties are heterogeneous, with a wide range of rise and decline times (e.g., Angus et al. 2019). Qualitatively similar “double peaked” light curves are seen in a fraction $\sim 3-9\%$ of ordinary-luminosity stripped-envelope supernovae (e.g., Das et al. 2023 and references therein), which however are frequently attributed to shock interaction between the supernova ejecta and CSM from pre-explosion mass-loss.
@@
It may not be a coincidence that the timescales of the observed bumps coincide both with the first light from the jet breakout [Eq. (1)] and the engine durations needed to power SLSNe emission. The luminosities of the bumps can also be accommodated within the energy budget of the engine (light blue cloud in Fig. 1). Kasen et al. (2016) proposed this early emission phase is powered by ejecta re-heating due to a delayed spherical shock breakout through the ejecta, distinct from the earlier supernova shock breakout and driven by the pressure of a pulsar-like nebula inflated inside the ejecta by the magnetar wind (see also Chen et al. 2016; Suzuki & Maeda 2016).
@@
In this Letter, we investigate whether this early emission phase can instead be powered by a collimated jet--from an engine with similar properties needed to power the full SLSNe light curve at later times--that successfully breaks out of the supernova ejecta. In §2.1, we establish a numerical simulation to track the jet propagation inside and outside the ejecta over a few weeks. In §2.2, we outline the post-processing calculations to estimate the resulting emission. In §3, we present our findings, demonstrating their consistency with observations of the early bumps in SLSNe-I. We discuss some implications of our results in §4.

### v2

# Late Jets, Early Sparks: Illuminating the Premaximum Bumps in Superluminous Supernovae

@@ Abstract
Superluminous supernovae (SLSNe) radiate $\gtrsim 10-100$ times more energy than ordinary stellar explosions, implicating a novel power source behind these enigmatic events. One frequently discussed source, particularly for hydrogen-poor (Type I) SLSNe, is a central engine such as a millisecond magnetar or accreting black hole. Both black hole and magnetar engines are expected to channel a fraction of their luminosity into a collimated relativistic jet. Using 3D relativistic hydrodynamical simulations, we explore the interaction of a relativistic jet, endowed with a luminosity $L_{\rm j}\approx 10^{45.5}\,{\rm erg\,s^{-1}}$ and duration $t_{\rm eng}\approx 10\,{\rm days}$ compatible with those needed to power SLSNe, launched into the envelope of the exploding star. The jet successfully breaks through the expanding ejecta, and its shocked cocoon powers ultraviolet/optical emission lasting several days after the explosion and reaching a peak luminosity $\gtrsim 10^{44}\,{\rm erg\,s^{-1}}$, corresponding to a sizable fraction of $L_{\rm j}$. This high radiative efficiency is the result of the modest adiabatic losses the cocoon experiences owing to the low optical depths of the enlarged ejecta at these late times, e.g., compared to the more compact stars in gamma-ray bursts. The luminosity and temperature of the cocoon emission match those of the “bumps” in SLSN light curves observed weeks prior to the optical maximum in many SLSNe. Confirmation of jet breakout signatures by future observations (e.g., days-long to weeks-long internal X-ray emission from the jet for on-axis observers, spectroscopy confirming large photosphere velocities $v/c\gtrsim 0.1$, or detection of a radio afterglow) would offer strong evidence for central engines powering SLSNe.
@@ 1 Introduction
The collapse of a rotating massive star giving birth to a magnetized spinning compact object – either a Kerr black hole or a neutron star – offers a compelling scenario for generating some of the most energetic transients in the Universe, including gamma-ray bursts (GRBs) and superluminous supernovae (SLSNe). While most GRBs with long durations are accompanied by stripped-envelope SNe with abnormally large kinetic energies (“broad-lined” Type Ibc SNe; e.g., Stanek et al. 2003; Woosley & Bloom 2006; Liu & Modjaz 2016), SLSNe are instead distinguished by their large radiated energies, which can exceed those of ordinary stellar explosions by a factor of $\gtrsim 10-100$ (e.g., Gal-Yam 2018; Nicholl 2021 for reviews). Type I (i.e., hydrogen-poor) SLSNe, in particular, may be associated with the birth of a millisecond (ms) neutron star with a strong magnetic field (“ms magnetar”; e.g., Kasen & Bildsten 2010; Woosley 2010; Nicholl et al. 2017; Vurm & Metzger 2021). The magnetized wind that accompanies the first tens of seconds in the life of a ms magnetar can feed a relativistic jet with the necessary luminosity and baryon loading (Lorentz factor) to power a GRB (e.g., Wheeler et al. 2000; Thompson et al. 2004; Bucciantini et al. 2008; Metzger et al. 2011), though the accretion-powered jet of a Kerr black hole remains a more commonly invoked scenario (e.g., Woosley 1993; MacFadyen & Woosley 1999; Gottlieb et al. 2023).
@@
What distinguishes the central engines of GRBs versus those that power SLSNe? Although the total energetics of both systems can be comparable, their main distinction is the duration of the central engine activity (Metzger et al., 2015). While GRB engines typically emit a sizable fraction of their energy (Panaitescu & Kumar, 2002) over seconds to minutes (comparable to the duration of the gamma-ray emission; e.g., Bromberg et al. 2012), the engines of SLSNe must instead heat the SN ejecta continuously from within over much longer timescales of days to weeks. This heating being substantially delayed after the time of the stellar explosion is crucial to explaining the high radiative output of SLSNe. A delay ensures that less deposited energy is “wasted” through adiabatic losses by occurring at later times after the optical depth through the expanding ejecta has dropped. Fig. 1 shows estimates of the engine duration and luminosity required to explain the rise times and peak luminosities of a large sample of Type I SLSNe (Gomez et al., 2024).
@@
In magnetar scenarios, the engine duration is typically related to the magnetic dipole spin-down timescale $t_{\rm sd}\propto B^{2}$ which varies considerably with the surface magnetic field strength ($B\sim 10^{14}$ G to power SLSNe vs. $B\gtrsim 3\times 10^{15}$ G to power a typical long GRB). Even at fixed $B$, differences in the magnetar’s mass accretion rate (Metzger et al., 2018) or the inclination angle of the magnetic dipole relative to the spin axis (Margalit et al., 2018) can lead to large differences in the jet power. So-called “ultralong” GRBs (e.g., Levan et al. 2014) require engine durations of hours, commensurate with their longer gamma-ray emission; these events may represent an intermediate case in which the SN’s luminosity is boosted only modestly above the baseline level set by 56Ni decay (Greiner et al., 2015; Metzger et al., 2015), the latter dominating in ordinary GRB SNe (e.g., Woosley et al. 2002).
@@
While a variety of indirect evidence supports a compact object residing at the center of at least some Type I SLSNe (e.g., Lunnan et al. 2015; Eftekhari et al. 2019; Nicholl et al. 2019; Vurm & Metzger 2021; Omand & Jerkstrand 2023), a central engine is not the only plausible power source for these events. A commonly invoked alternative is shock interaction between the fast SN ejecta and a compact wind or other circumstellar medium (CSM) that surrounds the progenitor star at the time of explosion (e.g., Chevalier & Irwin 2011; Chatzopoulos et al. 2013; Lunnan et al. 2018; Fraser 2020; Aguilera-Dena et al. 2020; Stevance & Eldridge 2021). Given the challenges in cleanly distinguishing engine from CSM interaction models, alternative tests of the engine hypothesis are sorely needed. Gamma-ray emission from a magnetar nebula on timescales of months to years after the ejecta becomes optically-thin offers one compelling test (e.g., Murase et al. 2015; Vurm & Metzger 2021; Murase et al. 2021); however, the low predicted gamma-ray flux at these late times makes detection challenging in all but the closest explosions (Acharyya et al., 2023). A more promising test is one that probes the central engine at earlier times closer to its maximal power.
@@
If a rotating compact object powers both GRBs and SLSNe, it is natural to ask whether SLSNe should also be accompanied by ultrarelativistic jets (see Soker, 2022, for a review of the role of jets in CCSNe). While the luminosities of SLSN engines are lower than those of GRB jets, the density of the external medium through which a putative SLSN jet must propagate is also lower. This is due to the larger ejecta radius $R_{\rm ej}\approx\,{v_{\rm ej}}t\sim 10^{13}-10^{14}$ cm at such late times tsimilar-$t\sim$ days--weeks after the explosion. Here, $\,{v_{\rm ej}}\approx 5000$ km s-1 is the speed of the SN ejecta, which for typical stripped-star progenitor radii $R_{\star}\sim 10^{11}$ cm, achieves homologous expansion on a timescale $\gtrsim R_{\star}/\,{v_{\rm ej}}\sim 10^{2}$ s. Analytic estimates show that even jet luminosities $\,{L_{\rm j}}\simeq 10^{45}$ erg s-1, much lower than for GRBs but comparable to those required of many Type I SLSNe engines (Fig. 1), are sufficient for jet breakout through the homologously expanding stellar ejecta (e.g., Quataert & Kasen 2012; Margalit et al. 2018; Gottlieb & Nakar 2022). In particular, for a jet of half-opening angle $\theta_{j}$ and luminosity $\,{L_{\rm j}}$, breakout can occur on a timescale (Eq. 45 in Gottlieb & Nakar, 2022):
$t_{\rm bo}\approx 5\,\frac{E_{\rm KE}}{3\times 10^{51}\,\,{{\rm erg}}}\frac{10^{45}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}}{\,{L_{\rm j}}}\left(\frac{\theta_{j}}{10^{\circ}}\right)^{4}\,\,{{\rm d}}\,,$ (1)
@@
where the prefactor assumes a non-magnetized jet propagation, as expected for extended-duration weak jets, and other quantities including the ejecta kinetic energy $E_{\rm KE}$ are normalized to characteristic values (Fig. 1). The jet breakout times of days indicated by Eq. (1) are notably similar to the typical engine durations (e.g., magnetar spin-down time) required to power the bulk of the SLSN emission.
@@
Figure 1: Required engine peak luminosity, $L_{\rm eng}$ and duration at this luminosity, $t_{\rm eng}$ (e.g., magnetar spin-down luminosity and timescale, respectively), to explain the observed peak luminosities, $L_{\rm pk}$, and rise times, $t_{\rm pk}$, of a large sample of Type I SLSN light curves (Gomez et al., 2024), for two different assumptions about the ejecta mass, $M_{\rm ej}=10M_{\odot}$ (black stars) and $3M_{\odot}$ (blue squares), bracketing the typical range inferred for Type I SLSNe (Nicholl et al., 2017). For comparison, we show the approximate range of timescales and luminosities of the early premaximum bumps observed in the light curves of Type I SLSNe (shaded blue ellipse; see e.g. Nicholl & Smartt 2016). A multiday-lasting jet from the same central engine responsible for powering the bulk of the SLSNe is thus capable of powering the early bumps, assuming that the jet luminosity, $\,{L_{\rm j}}$, is a fraction $\sim 1-10\%$ of $L_{\rm eng}$. The assumed jet properties for the numerical simulation presented in this paper are illustrated with an orange square, while bidirectional arrows indicate the degeneracy in engine properties $L_{\rm eng}\propto t_{\rm eng}^{-1}$ that result in the same hydrodynamic evolution for fixed ejecta properties. Details: For each event, we estimate the ejecta kinetic energy, $E_{\rm KE}=M_{\rm ej}\,{v_{\rm ej}}^{2}/2$, by equating the observed SN rise-timescale to the diffusion timescale for a centrally-concentrated energy source, $t_{\rm diff}\simeq 0.4(\kappa M_{\rm ej}/\,{v_{\rm ej}}c)^{1/2}$ (Khatami & Kasen, 2019). Assuming that the explosion energy $E_{\rm KE}$ comes exclusively from the engine, i.e. $E_{\rm eng}=L_{\rm eng}t_{\rm eng}=E_{\rm KE}$, we estimate the required engine timescale assuming that the engine luminosity decays similarly to magnetic-dipole spin-down $L_{\rm pk}=(E_{\rm eng}/t_{\rm eng})(1+t_{\rm pk}/t_{\rm eng})^{-2}$, for $t_{\rm eng}<t_{\rm pk}$. If the engine contributes only a fraction of the explosion energy $E_{\rm KE}$, then a greater engine duration is needed to explain the observed luminosities and points will shift towards the lower right part of the diagram as indicated by the arrow.
@@
Do observations rule out jets in SLSNe? With one interesting exception (Levan et al. 2013), no X-ray emission has been detected to faint limits from a sample of 26 Type I SLSNe on timescales of months to years after the explosion (e.g., Margutti et al. 2018). However, this is not necessarily constraining because the spectral energy distribution of any (internal) jetted emission is uncertain; and even if the jet efficiently powers X-rays, they may only be detectable at early times and for observers within the narrow opening angle of the jet. In principle, radio afterglow emission from the jet as it interacts with surrounding CSM may be detectable at late times even for slightly off-axis observers; however, again (with one or two interesting exceptions; Eftekhari et al. 2019; Margutti et al. 2023111In PTF10hgi (Eftekhari et al., 2019; Mondal et al., 2020), radio emission detected starting 7.5 years after the explosion may originate from a magnetar nebula (e.g., Margalit & Metzger 2018; Omand et al. 2018), while the late radio emission in SN 2017ens (Margutti et al., 2023) originates from the SN ejecta colliding with CSM.) no radio detections of Type I SLSNe have been made (e.g., Coppejans et al. 2018; Eftekhari et al. 2021; Hatsukade et al. 2021). As we shall discuss, the resulting constraints on the jet power are sensitive to the observer viewing angle, CSM density, and uncertain microphysical parameters of the shock.
@@
A cleaner signature of a successful jet in Type I SLSNe may come at earlier times, on timescales comparable to the engine duration of days to weeks (Margalit et al., 2018). In ordinary GRB jets, the compact size of the shocked stellar debris (the so-called “cocoon”; e.g., Gottlieb et al. 2018a) results in most of the cocoon radiation being initially trapped and only subsequently released after many expansion times. By contrast, for the more sustained jet-ejecta interaction expected for longer-lived SLSN jets interacting with substantially larger ejecta, the lower optical depths (smaller trapping radius in the ejecta frame) will alter the observed cocoon light curve, allowing it to track energy input from the engine more faithfully.
@@
A suggestive hint comes from the early maxima (“bumps”) of luminosity $\sim 10^{43}-10^{45}$ erg s-1 observed in some Type I SLSNe optical light curves on timescales of days to a week after the explosion, well before the main SN peak (Leloudas et al., 2012; Nicholl et al., 2015). Although such premaximum bumps are not detected in all events due to observational limitations, Nicholl & Smartt (2016) show that they are likely common features in SLSNe. Subsequent studies have shown that early maxima are not present in all SLSNe, and that their properties are heterogeneous, with a wide range of rise and decline times (e.g., Angus et al. 2019). Qualitatively similar “double peaked” light curves are seen in a fraction $\sim 3-9\%$ of ordinary-luminosity stripped-envelope SNe (e.g., Das et al. 2023 and references therein), which however are frequently attributed to shock interaction between the SN ejecta and CSM from pre-explosion mass-loss.
@@
It may not be a coincidence that the timescales of the observed bumps coincide both with the first light from the jet breakout [Eq. (1)] and the engine durations needed to power SLSN emission. The luminosities of the bumps can also be accommodated within the energy budget of the engine (light blue cloud in Fig. 1). Kasen et al. (2016) proposed that this early emission phase is powered by ejecta reheating due to a delayed spherical shock breakout through the ejecta, distinct from the earlier SN shock breakout and driven by the pressure of a pulsar-like nebula inflated inside the ejecta by the magnetar wind (see also Chen et al. 2016; Suzuki & Maeda 2016; Moriya et al. 2022).
@@
In this Letter, we investigate whether this early emission phase can instead be powered by a collimated jet--from an engine with similar properties needed to power the full SLSN light curve at later times--that successfully breaks out of the SN ejecta. In §2.1, we establish a 3D numerical simulation to track the jet propagation inside and outside the ejecta over several weeks (see also Ramirez-Ruiz & MacFadyen, 2010; De Colle et al., 2022, for 2D simulations of jet-SN ejecta interaction). In §2.2, we outline the post-processing calculations to estimate the resulting emission. We present our findings in §3, demonstrating their consistency with observations of the early bumps in Type I SLSNe. We discuss some implications of our results in §4.